If last year, 47% of internet users wanted to decrease the time spent
online and this year 41 of the 70 persons surveyed wants to, at the
5% significance level is there evidence to support a change in the
rate?

Answer by ewatrrr(24785)

Ho: p = 0.47 (last year)
Ha: p ≠ .47
-----------------------------------------
n = 70
sample proportion = p-hat = 41/70 = 0.586
---
test stat:: z(0.586) = (0.586-0.47)/sqrt[0.47*0.53/70] = .116/.0035 = 32.59
critical value: ±1.96
Reject Ho: 32.59 > 1.96
at the 5% significance level, the evidence supports a change in the rate